How many milliliters of water must be added to 12 ml of a 60% salt solution in order to produce a 42% salt solution? Round your answer to the nearest tenth.
Answer by stanbon(75887)

How many milliliters of water must be added to 12 ml of a 60% salt solution in order to produce a 42% salt solution? Round your answer to the nearest tenth.
-------
Equation:
salt + salt = salt
0*x + 0.60*12 = 0.42(x+12)
Multiply thru by 100 to get:
0 + 60*12 = 42x + 42*12
42x = 18*12
x = 5.14 ml (amt. of water needed)
=====================================
Cheers,
Stan H.
==========